the sum of the ages of tina and fred is 38 years. four years ago,the age of tina was the square of freds age. find their ages
Answer by Alan3354(69443)

the sum of the ages of tina and fred is 38 years. four years ago,the age of tina was the square of freds age. find their ages
----------
4 years ago the sum was 30
T + F = 30
T = F^2
F^2 + F - 30 = 0
F = 5
T = 25
--------------
--> F = 9 now
T = 29 now